OpenStudy (jackfrostt3482):

Help Please asap, I will give medal and fan. I will post equation below

OpenStudy (sebastianwininger):

Ok i help

OpenStudy (jackfrostt3482):

ok, hold on :)

OpenStudy (sebastianwininger):

alright then

OpenStudy (jackfrostt3482):

What is the volume of a cone with a radius of 18 cm and a height of 10 cm? Use 3.14 for pi. Round your answer to the nearest tenth.

OpenStudy (3mar):

I hope that helps! Salam!

OpenStudy (jackfrostt3482):

first what its \[\frac{1}{3} \times \pi\]

OpenStudy (jackfrostt3482):

I think i got the answer to that earlier but i cant remember and i put the answer into the equation and it was wrong...

OpenStudy (jskhupmang):

r=18, h=10, pi=3.14 plug them into the formula: 1/3 pi r^2 h=1/3(3.14)(18)^2(10)= 3391.1999.. but round to the tenth so 3391.2 is the ans.

OpenStudy (jskhupmang):

@jackfrostt3482 isn't it correct?

OpenStudy (jackfrostt3482):

I think so

OpenStudy (jackfrostt3482):

but how did you \[\frac{1}{3}\times \pi\]

OpenStudy (jackfrostt3482):

or what did you get....

OpenStudy (jskhupmang):

1/3 = 0.333333.... never ending and multiply with 3.14 which is the pi.

OpenStudy (jskhupmang):

0.33333... times 3.14 = 1.0466666.... got it?

OpenStudy (jackfrostt3482):

yeah